OpenStudy (anonymous):

The results of surveying 100 residents of a city reported that 40 read the daily morning paper, 70 read the daily evening paper & 20 read neither. How many: 1. Read at least 1 paper? 2. Read both? 3. Read exactly one daily paper?

OpenStudy (anonymous):

1. 100 - 20 = 80 2. 100 - (100 - 70) - (100 - 40) = 10 3. 80 - 10 = 70

OpenStudy (anonymous):

thanks but for my answer key its 2) 30 c)50

hero (hero):

If you have to post another question, that means you didn't master it yet.

OpenStudy (anonymous):

i didn't master these problems?

OpenStudy (anonymous):

ya i have the answer for it...i would just like to check it and discuss it

hero (hero):

Maybe @apoorvk thought you did

OpenStudy (anonymous):

no this is another question

hero (hero):

They're similar questions, however

hero (hero):

You have three independent events and you have to find probabilities based on these events.

OpenStudy (anonymous):

ya its all probability and venn diagrams...i know more or less how to get it but my problem is using the correct formula to get it

OpenStudy (anonymous):

for instance this is how i would get it (but not sure if theres a better formula)

hero (hero):

Well, using a different setup might help.

OpenStudy (anonymous):

well thats why i am here to ask for suggestions

OpenStudy (anonymous):

for instance it says 100 respondents were surveyed and 20 ppl said they read none so 100-20= 80 so 80 have to read at least ONE

OpenStudy (anonymous):

for both: well i see that 40 ppl raed morning and 70 ppl read evening so thats a total of 110 distributed newspaper so 110 newspaper-80 (who read at least one) = 30?

OpenStudy (anonymous):

so theres an extra 30 newspaper left, which means 30 must read an extra one

OpenStudy (anonymous):

so for reading exactly one 70-30= 40 evening 40-10= 10 morning like in my head i can see how it plays out...but i struggle trying to find the correct formu

OpenStudy (apoorvk):

Well well.. let us call Mr. Venn, and draw a Venn Diagram here Hmm so let S be the sample space of the 100 people being surveyed. Region A= represents no. of morning readers, and region B = the no. of evening readers So, nS=100 nA=40 nB=70 Now ofcourse the total no. of people who read any or atleast one newspaper is 100-20, since 20 don't read any. so, no. of readers (atleast one) =80 And, if you notice no. of readers is the union of setA and setB, since it includes morning and evening readers both. so, n(AUB) = 80 And (AintersectionB) is the set of people who read BOTH newspapers. now, if you have studied set theory, n(AUB) = nA + nB - n(AintersectionB) so, n(AintersectionB) = 70 + 40 - 80 = 30 (In case you don't understand that, try to think analytically, or else I 'll explain it to you) And so, if 30 people read both newspapers, the rest of the readers, i.e. 80-30 = 50 read just one of the two newspapers.